when two is multiplied by one of the following number. the resulting number is reverse the number when two added to it. what is that number.
(19,33,47,51)

Answer by fractalier(6550)

I did this by trial and error.
19 * 2 = 38 does not equal 19 + 2 reversed
33 * 2 = 66 does not equal 33 + 2 reversed
51 * 2 = 102 does not equal 51 + 2 reversed
but 47 * 2 = 49 does equal 47 + 2 reversed
47 is the answer.